# Differentiable Weightless Neural Networks

### v1

@@ 4.1 DWNs on FPGAs
Table 1: Implementation results for DWNs and prior efficient inference models on a Z-7045 FPGA. Model could not be synthesized; hardware values are approximate. ULEEN used augmented data for MNIST; we present MNIST results with and without augmentation.
@@ 4.2 DWNs on Microcontrollers
We use two strategies for implementing DWNs on the Nano. Our first approach uses aggressive bit packing to minimize memory usage, allowing us to fit more complex models on the device. For instance, the 64 entries of a LUT-6 can be packed in 8 bytes of memory, and the six indices for its inputs can be stored in 7.5 bytes by using 10-bit addresses (for more details on our bit-packing strategy, see Appendix D). However, this approach needs to perform bit manipulation to unpack data, which is fairly slow. Therefore, we also explore an implementation without bit-packing, which greatly increases inference speed but reduces the maximum size (and therefore accuracy) of feasible models.
@@ 4.3 DWNs for Ultra-Low-Cost Chips
To assess the viability of DWNs for ultra-low-cost chip implementations, we analyze their performance in terms of accuracy and NAND2 equivalent circuit area, comparing them with Tiny Classifiers (Iordanou et al., 2024), a SOTA work for ultra-low-cost small models, and DiffLogicNet. The datasets for this analysis are those shared between the Tiny Classifiers (see Appendix G) and AutoGluon (to be used in the next subsection) studies, providing a consistent basis for comparison. We also adhere to their data-splitting methods, using 80% of the data for the training set and 20% for the testing set.
Our DWN, utilizing the Learnable Reduction technique with LUT-2s, is designed to inherently learn two input binary logic, which directly correlates to logic gate formation in a logic circuit. The NAND2 equivalent size of our model is calculated by converting each LUT-2 into its NAND2 equivalent (e.g., a LUT-2 representing an OR operation equates to 3 NAND gates). For DiffLogicNet, we adopt a similar approach, translating the converged binary logic nodes into their NAND2 equivalents, plus the additional NAND2 equivalent size required for each class output popcount (Appendix E), as per their model architecture. Notably, our DWN model, due to Learnable Reduction, does not incur this additional computational cost. For Tiny Classifiers we utilize the results reported in their paper.
@@ 4.4 DWNs on Tabular Data
For training our DWN, we adopted an AutoML-like approach: setting the number of layers to 3 for small datasets and 5 for larger ones. The number of LUTs per layer was adjusted to align with the final model sizes of XGBoost, thereby maintaining a comparable model size. See Appendix I for more model configuration details.
@@ Appendix C DiffLogicNet’s Binary Logic Node
Appendix D Bit-Packed Microcontroller Implementation Details
@@
Appendix E Estimating NAND2 Equivalents for Popcount Circuits
@@
Appendix F Expanded Details on Tabular Data Comparisons for Tiny Models
@@
Appendix G List of Datasets Used
@@
Appendix H Ablation Studies
See Tables 9, 10, 11 and 12.
@@ Appendix H Ablation Studies
Appendix I Model Configurations
See Tables 13, 14, 15 and 16.

### v2

@@ 4.1 DWNs on FPGAs
Table 1: Implementation results for DWNs and prior efficient inference models on a Z-7045 FPGA with input buffers and decompression implemented, with data being passed to the board using its I/O pins, which have a bandwidth of 112 bits per cycle, and the clock is limited to 200 MHz for a fair comparison to prior work. For results in out-of-context mode, where data is assumed to already be on the board and the clock is set to 250 MHz, see Appendix D. Model could not be synthesized; hardware values are approximate. ULEEN used augmented data for MNIST; we present MNIST results with and without augmentation.
@@ 4.2 DWNs on Microcontrollers
We use two strategies for implementing DWNs on the Nano. Our first approach uses aggressive bit packing to minimize memory usage, allowing us to fit more complex models on the device. For instance, the 64 entries of a LUT-6 can be packed in 8 bytes of memory, and the six indices for its inputs can be stored in 7.5 bytes by using 10-bit addresses (for more details on our bit-packing strategy, see Appendix E). However, this approach needs to perform bit manipulation to unpack data, which is fairly slow. Therefore, we also explore an implementation without bit-packing, which greatly increases inference speed but reduces the maximum size (and therefore accuracy) of feasible models.
@@ 4.3 DWNs for Ultra-Low-Cost Chips
To assess the viability of DWNs for ultra-low-cost chip implementations, we analyze their performance in terms of accuracy and NAND2 equivalent circuit area, comparing them with Tiny Classifiers (Iordanou et al., 2024), a SOTA work for ultra-low-cost small models, and DiffLogicNet. The datasets for this analysis are those shared between the Tiny Classifiers (see Appendix H) and AutoGluon (to be used in the next subsection) studies, providing a consistent basis for comparison. We also adhere to their data-splitting methods, using 80% of the data for the training set and 20% for the testing set.
Our DWN, utilizing the Learnable Reduction technique with LUT-2s, is designed to inherently learn two input binary logic, which directly correlates to logic gate formation in a logic circuit. The NAND2 equivalent size of our model is calculated by converting each LUT-2 into its NAND2 equivalent (e.g., a LUT-2 representing an OR operation equates to 3 NAND gates). For DiffLogicNet, we adopt a similar approach, translating the converged binary logic nodes into their NAND2 equivalents, plus the additional NAND2 equivalent size required for each class output popcount (Appendix F), as per their model architecture. Notably, our DWN model, due to Learnable Reduction, does not incur this additional computational cost. For Tiny Classifiers we utilize the results reported in their paper.
@@ 4.4 DWNs on Tabular Data
For training our DWN, we adopted an AutoML-like approach: setting the number of layers to 3 for small datasets and 5 for larger ones. The number of LUTs per layer was adjusted to align with the final model sizes of XGBoost, thereby maintaining a comparable model size. See Appendix J for more model configuration details.
@@ Appendix C DiffLogicNet’s Binary Logic Node
Appendix D Out-of-Context Results
In our primary Table 1, we reported FPGA results where input buffers and decompression were implemented, assuming data is streamed to the board using I/O pins (112 bits per cycle on the Zynq Z-7045 FPGA) and limited the clock speed to 200 MHz. These design choices were made to enable a fair comparison to prior work, particularly FINN, which adheres to similar constraints.
However, we noticed that other papers report results assuming data is already available on the board without accounting for streaming or decompression overheads (i.e., operating in out-of-context mode). To provide a comprehensive comparison, we present these out-of-context results in the Table 7, following the methodology of other papers by setting the clock speed to the board’s maximum of 250 MHz.
In this configuration, our model achieves an effective throughput of 250 million images per second. This is because our design is fully pipelined, taking only 4 ns per image during inference as the model is compact enough to process one image per clock cycle.
Table 7: Implementation results for DWNs on a Z-7045 FPGA compiled in out-of-context mode with the clock set to an optimal 250 MHz (4 ns per cycle), the maximum frequency of the board.
Dataset Model Test AccuracyParameter Size (KiB) Time/Sample (ns) Throughput (Samples/s) LUTs (1000s)
MNIST DWN ($n$=2; lg) 98.27 5.9 4 250 M 6.5
DWN ($n$=6; sm) 97.80 11.7 4 250 M 2.1
DWN ($n$=6; lg) 98.31 23.4 4 250 M 4.1
DWN ($n$=6; lg; +aug) 98.77 23.4 4 250 M 4.1
FashionMNIST DWN ($n$=2) 89.12 7.8 4 250 M 8.1
DWN ($n$=6) 89.01 31.3 4 250 M 6.2
KWS DWN ($n$=2) 70.92 2.9 4 250 M 3.3
DWN ($n$=6) 71.52 12.5 4 250 M 3.3
ToyADMOS/car DWN ($n$=2; sm) 86.68 0.9 4 250 M 1.0
DWN ($n$=6; sm) 86.93 3.1 4 250 M 0.8
DWN ($n$=6; lg) 89.03 28.1 4 250 M 5.5
CIFAR-10 DWN ($n$=2) 57.51 23.4 4 250 M 45.7
DWN ($n$=6) 57.42 62.5 4 250 M 16.7
Appendix E Bit-Packed Microcontroller Implementation Details
@@
Appendix F Estimating NAND2 Equivalents for Popcount Circuits
@@
Appendix G Expanded Details on Tabular Data Comparisons for Tiny Models
@@
Appendix H List of Datasets Used
@@
Appendix I Ablation Studies
See Tables 10, 11, 12 and 13.
@@ Appendix I Ablation Studies
Appendix J Model Configurations
See Tables 14, 15, 16 and 17.